# Accurate quantum-centric simulations of supramolecular interactions

### v1

@@ III Results
Figure 7: The total energy of a methane dimer predicted with SQD (16e,24o) at 3.638 Å distance between the monomers as the function of $1/d^{\prime}\cdot 10^{6}$. Dashed green line with triangular markers shows SQD (16e,24o) results. Horizontal dashed cyan line indicate the total energy from HCI (16e,24o) calculation.
We show the decrease in the SQD (16e,24o) total energy for the methane dimer at 3.638 Å with the increase of $|\widetilde{\chi}_{b}|$ from $5.0\cdot 10^{3}$ to $8.5\cdot 10^{3}$ in Fig. 7. The differences between the total energies predicted with SQD (16e,24o) and HCI (16e,24o) reduce from 40.8 milliHartree to 23.4 milliHartree when the $|\widetilde{\chi}_{s}|$ is increased from $5.0\cdot 10^{3}$ to $8.5\cdot 10^{3}$. This increase in the number of samples corresponds to an decrease of $1/d^{\prime}\cdot 10^{6}$ from 1.7 to 0.7. Here $d^{\prime}$ is the number of the configurations with the contributions to the wavefunction that are higher than the threshold of $1.0\cdot 10^{-8}$. We believe that a further increase in the number of samples will allow us to advance the accuracy of SQD (16e,24o) calculations. To make SQD (16e,24o) calculations of the methane dimer more computationally feasible we are currently exploring parallelization options for calculations on this system as well as the analysis of the configurations with low contributions to the total energies.
@@ IV Conclusions and Outlook
Our findings demonstrate that SQD is capable of capturing noncovalent interactions between molecules at the level of theory chosen, with potential energy surfaces that closely align with those obtained through classical computational methods. We examine the binding energies of the water and methane dimers by comparing SQD with an analogous classical method, namely HCI. We also compare SQD against the CCSD(T) method, which is considered the gold standard for calculations of binding energies [carsky2010recent]. This comparison aims to evaluate the accuracy of SQD and to understand how the nature of the PES changes with different active-space selections. The ability of HCI and SQD to recover the dispersion interaction is highly dependent on the size and nature of the active space, which is especially critical for predicting the binding energy of the methane dimer. In fact, a previous study by Hapka et al.[hapka2020dispersion] demonstrated that the ability of supramolecular multiconfigurational interaction calculations to recover the dispersion energy depends on the size of the active space and can be improved with systematic expansion of the active space.
The results obtained here demonstrate the improvements both in terms of accuracy and scale of quantum computations on chemical problems, enabling, on current quantum processors, use cases previously thought to belong to the fault-tolerant domain, such as the largest active space considered here for methane, which has  1.3M Pauli operators. Further examples of problems that coule be enabled by our approach include quantum computing simulations of chemical reactivity of $CO_{2}$-fixating ruthenium catalyst proposed by Burg et al. [von2021quantum], Ibrutinib drug simulations proposed by Blunt et al.[blunt2022], and the drug-discovery workflows proposed by Pyrkov et al.[PYRKOV2023103675] and Kumar et al.[10466774], as well as multiple stages of drug optimization as described by Bonde et al.[Bondechapter7] The SQD method allows for simulations of systems with qubit counts that is essential for projection-based embedding algorithm proposed by Ralli et al.[Ralli2024] and can enhance the viability of fragment-based quantum computing simulations. Previously, VQE-based fragment molecular orbital (FMO),[Lim2024] divide and conquer (DC),[Yoshikawa2022] and density matrix embedding theory (DMET)[kawashima2021optimizing, iijima2023accurate] simulations were limited to very simple illustrative systems. Shang et al. proposed a DMET-based massively parallel quantum computing approach based on VQE, but execution of their methodology was only possible on a quantum simulator rather than actual hardware.[Shang2023]

### v2

@@ III Results
Figure 7: Extrapolation of SQD (16e,24o) total energies for the methane dimer at 3.638 Å distance between the carbon atoms. Extrapolations are done using four points with $|\tilde{\chi}_{b}|$ of $5.5\cdot 10^{3}$, $6.5\cdot 10^{3}$, $7.5\cdot 10^{3}$, and $8.5\cdot 10^{3}$. Hamiltonian variance ($\Delta H$) is calculated as $\Delta H=\langle\psi^{(k)}|\hat{H}^{2}|\psi^{(k)}\rangle-\langle\psi^{(k)}|\hat{H}|\psi^{(k)}\rangle^{2}$. Green triangles indicate SQD (16e,24o) energies. Grey dashed lines indicate extrapolated SQD (16e,24o) energies. Dashed red line indicates HCI (16e,24o) energies. Error bars indicate magnitude of error estimate in extrapolation.
We show the decrease in the SQD (16e,24o) total energy for the methane dimer at 3.638 Å with the increase of $|\widetilde{\chi}_{b}|$ from $5.0\cdot 10^{3}$ to $8.5\cdot 10^{3}$ in Fig. 7. The differences between the total energies predicted with SQD (16e,24o) and HCI (16e,24o) reduce from 31.8 milliHartree to 25.2 milliHartree when the $|\widetilde{\chi}_{s}|$ is increased from $5.0\cdot 10^{3}$ to $8.5\cdot 10^{3}$. The extrapolated total energy based on SQD (16e,24o) simulations with $|\tilde{\chi}_{b}|$ of $5.5\cdot 10^{3}$, $6.5\cdot 10^{3}$, $7.5\cdot 10^{3}$, and $8.5\cdot 10^{3}$ agrees with HCI (16e,24o) results within 2.12 milliHartree. Magnitude of the error estimate in extrapolation is $\pm$ 3.10 milliHartree. We believe that a further increase in the number of samples will allow us to advance the accuracy of SQD (16e,24o) calculations. To make SQD (16e,24o) calculations of the methane dimer more computationally feasible we are currently exploring parallelization options for calculations on this system as well as the analysis of the configurations with low contributions to the total energies.
@@ IV Conclusions and Outlook
Our findings demonstrate that SQD is capable of capturing noncovalent interactions between molecules at the level of theory chosen, with potential energy surfaces that closely align with those obtained through classical computational methods. We examine the binding energies of the water and methane dimers by comparing SQD with an analogous classical method, namely HCI. We also compare SQD against the CCSD(T) method, which is considered the gold standard for calculations of binding energies [90]. This comparison aims to evaluate the accuracy of SQD and to understand how the nature of the PES changes with different active-space selections. The ability of HCI and SQD to recover the dispersion interaction is highly dependent on the size and nature of the active space, which is especially critical for predicting the binding energy of the methane dimer. In fact, a previous study by Hapka et al.[115] demonstrated that the ability of supramolecular multiconfigurational interaction calculations to recover the dispersion energy depends on the size of the active space and can be improved with systematic expansion of the active space.
The results obtained here demonstrate the improvements both in terms of accuracy and scale of quantum computations on chemical problems, enabling, on current quantum processors, use cases previously thought to belong to the fault-tolerant domain, such as the largest active space considered here for methane, which has  1.3M Pauli operators. Further examples of problems that coule be enabled by our approach include quantum computing simulations of chemical reactivity of $CO_{2}$-fixating ruthenium catalyst proposed by Burg et al. [116], Ibrutinib drug simulations proposed by Blunt et al.[117], and the drug-discovery workflows proposed by Pyrkov et al.[118] and Kumar et al.[119], as well as multiple stages of drug optimization as described by Bonde et al.[120] The SQD method allows for simulations of systems with qubit counts that is essential for projection-based embedding algorithm proposed by Ralli et al.[121] and can enhance the viability of fragment-based quantum computing simulations. Previously, VQE-based fragment molecular orbital (FMO),[122] divide and conquer (DC),[123] and density matrix embedding theory (DMET)[65, 124] simulations were limited to very simple illustrative systems. Shang et al. proposed a DMET-based massively parallel quantum computing approach based on VQE, but execution of their methodology was only possible on a quantum simulator rather than actual hardware.[125]
@@
Acknowledgements.
The authors gratefully acknowledge financial support from the National Science Foundation (NSF) through CSSI Frameworks Grant OAC-2209717 and from the National Institutes of Health (Grant Numbers GM130641).